OpenStudy (anonymous):

what is the rule for 1,4,9,16

OpenStudy (anonymous):

are you looking for the explicit or recursive answer?

OpenStudy (anonymous):

These numbers are jusr perfect square starting with 1. That is, 1^2 = 1 2^2 = 4 3^2 = 9, etc. So the rule is x^2, x is an integer and starting with 1.

OpenStudy (anonymous):

sorry i just wanted to say this is my most favorite rule EVER!

OpenStudy (anonymous):

wouldnt it really be n^2

OpenStudy (anonymous):

You can call it n^2, y^2, x^2, (anythi letter)^2...lol

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well the "correct" letter is n because when you display the sequence in summation notation the correct "equation" uses the variable n. ex: \[\sum_{n=1}^{4}n^2\]

OpenStudy (anonymous):

correct me if im wrong:)

OpenStudy (anonymous):

n is often used, but many texts use i, etc...

OpenStudy (anonymous):

ok my bad

OpenStudy (anonymous):

no i think n is the proper variable

OpenStudy (anonymous):

Then use n.

OpenStudy (anonymous):

and just for your information @Hiiiiiiiiiiiii n^2 would be an explicit equation